# Learning from Suboptimal Data in Continuous Control via Auto-Regressive Soft Q-Network

### v1

@@ B.4 Hyper-parameters
Table 1: Typical hyper-parameters of ARSQ in D4RL and RLBench.
The hyperparameters of ARSQ are presented in Table 1. We provide the typical hyperparameters for ARSQ in D4RL (hopper-medium) and RLBench (Open Oven). In RLBench, ARSQ employs RandomShift (Yarats et al., 2021) for image augmentation. Additionally, ARSQ utilizes SiLU (Hendrycks & Gimpel, 2016) and LayerNorm (Ba, 2016) as activation functions in RLBench.
@@ C.1 Motivating Example Setup
  • Q function given by independent action decomposition is an example of DecQN (Seyde et al., 2022), as well as in CQN (Seo et al., 2024), which features just a single coarse-to-fine level. In this setting, we employ separate tabular Q functions, $Q(s,a_{1})$ and $Q(s,a_{2})$, for action dimension 1 and action dimension 2. The Q function is learned by gradient descent.
@@ C.2 Environment and Dataset
To better demonstrate the suboptimal nature of the datasets, we have plotted a histogram of the data chunk rewards, as shown in Fig. 10.
@@ C.3 Baselines and Evaluation Details
In RLBench, we use the baselines and results of CQN, DrQ-v2+, DrQ-v2, ACT, and CBC as reported in the original CQN paper (Seo et al., 2024).
@@ Appendix D Additional Results
Figure 14: Sensitivity of temperature coefficient $\alpha$ in D4RL and RLBench.
@@
Figure 15: D4RL results per task on different demonstration quality.
@@ Appendix E Computational Cost Analysis
Table 2: Computational time in D4RL and RLBench (ms).

### v2

@@ B.4 Hyper-parameters
Table 2: Typical hyper-parameters of ARSQ in D4RL and RLBench.
@@
The hyperparameters of ARSQ are presented in Table 2. We provide the typical hyperparameters for ARSQ in D4RL (hopper-medium) and RLBench (Open Oven). In RLBench, ARSQ employs RandomShift (Yarats et al., 2022) for image augmentation. Additionally, ARSQ utilizes SiLU (Hendrycks & Gimpel, 2016) and LayerNorm (Ba, 2016) as activation functions in RLBench.
@@ C.1 Motivating Example Setup
  • Q function given by independent action decomposition is an example of DecQN (Seyde et al., 2023), as well as in CQN (Seo et al., 2024), which features just a single coarse-to-fine level. In this setting, we employ separate tabular Q functions, $Q(s,a_{1})$ and $Q(s,a_{2})$, for action dimension 1 and action dimension 2. The Q function is learned by gradient descent.
@@ C.2 Environment and Dataset
To better demonstrate the suboptimal nature of the datasets, we plot a histogram of the data chunk rewards, as shown in Fig. 10.
@@ C.3 Baselines and Evaluation Details
In RLBench, we adopt DrQ-v2+, an optimized variant of DrQ-v2 proposed by (Seo et al., 2024), as our baseline. DrQ-v2+ incorporates several optimization strategies introduced in the CQN algorithm. Specifically, compared to DrQ-v2, DrQ-v2+ employs a distributional critic instead of a standard critic network, utilizes an exploration strategy with small Gaussian noise, and features optimized network architectures and hyperparameters tailored for RLBench tasks. These enhancements strengthen DrQ-v2+’s performance, making it a more robust baseline than DrQ-v2. Additionally, DrQ-v2+ has been open-sourced by (Seo et al., 2024).
@@ Appendix D Additional Results
Figure 14: Sensitivity of temperature coefficient $\alpha$, evaluated on hopper-medium from D4RL and Open Oven from RLBench over three random seeds.
@@
Training Curves of D4RL Main Results.
In Section 5.1, we discuss the converged performance of ARSQ, CQN, and BC. The training curves for each task are shown in Fig. 15. ARSQ converges after approximately 25,000 to 50,000 environment steps and generally outperforms the CQN and BC baselines across most tasks. This further demonstrates ARSQ’s strength in managing suboptimal data.
Figure 15: Training curves of D4RL main results, evaluated over three random seeds.
@@
Figure 16: D4RL results per task on different demonstration quality, evaluated over three random seeds.
@@ Appendix E Computational Cost Analysis
Appendix F Performance under Fully Online Setting
In addition to the main experimental results presented in Sec. 5.1, we assess the performance of ARSQ in a fully online setting. We compare the fully online reinforcement learning performance of ARSQ and CQN on the hopper task, using PPO(Schulman et al., 2017) as a baseline for comparison. The results are depicted in Fig. 18, with all experiments conducted using three random seeds. We also illustrate the performance of both vanilla and offline ARSQ using the one of the hopper dataset, specifically hopper-medium-replay.
Online ARSQ achieves a similar converged performance to vanilla ARSQ, albeit requiring more environment steps. This highlights the importance of using offline datasets to enhance sample efficiency. Furthermore, ARSQ with online interaction achieves a higher final performance than in the offline setting, suggesting the necessity of online interaction to enhance policy performance. Additionally, online ARSQ demonstrates greater sample efficiency than CQN and PPO, underscoring its potential as a versatile reinforcement learning algorithm.
Figure 18: Performance under fully online settings, on hopper task or hopper-medium-replay dataset.
Appendix G Error Analysis of Q Prediction and Action Discretization
To further examine the error introduced by action discretization, as discussed in Sec. 4, we designe a more complex case study similar to Fig. 1. We create a one-step environment featuring a two-dimensional action space $(a_{1},a_{2})\in\mathcal{A}=[-1,1]^{2}\subset\mathbb{R}^{2}$. The agent performs an action at the initial timestamp, receives a reward, and the episode concludes. The Q function’s ground truth landscape, akin to the reward landscape, is illustrated in Fig. 19(a). There is one optimal action mode, two sub-optimal action modes, and two negative action modes.
We uniformly sample 2,000 data points from the environment to form a dataset. This dataset is then used to train agents with independent Q decomposition for each action dimension, ARSQ without hierarchical coarse-to-fine action discretization, and the standard ARSQ. The resulting Q landscapes are displayed in Fig. 19, and Q prediction errors are displayed in Fig. 20. When independently decomposing Q for each action dimension, the agent learns a blurred Q landscape, complicating the identification of optimal actions. ARSQ without coarse-to-fine action discretization produces a Q landscape similar to the vanilla ARSQ but with more ”glitches,” likely because too many action bins make it difficult for the dataset to cover them comprehensively. This underscores the importance of coarse-to-fine action discretization.
Furthermore, we sample 1,000 data points in the proposed environment and calculate the Q prediction error against the ground truth for all three methods discussed, over three random seeds. The results are presented in Tab. 4. Independent Q decomposition results in a significant error increase compared to ARSQ and continuous Q learning. Additionally, ARSQ without coarse-to-fine action discretization also results in higher Q error, further highlighting the necessity of our action discretization strategy.
(a) Ground truth.
(b) Independent action decomposition.
(c) ARSQ w/o coarse-to-fine discretization.
(d) ARSQ.
Figure 19: Visualization of Q prediction with different action discretization strategy.
(a) Independent action decomposition.
(b) ARSQ w/o coarse-to-fine discretization.
(c) ARSQ.
Figure 20: Q prediction errors with different action discretization strategy.
Table 4: Q prediction error with different action discretization strategy.
Discretization Method Error
Independent Decomposition 17.57 ± 0.67
ARSQ w/o Coarse-to-fine 0.50 ± 0.21
ARSQ 0.16 ± 0.02